OpenStudy (anonymous):

Simplify. Square root 3/25 plz help

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{5}\]

OpenStudy (anonymous):

\[\sqrt{\frac{3}{25}}\] \[\frac{\sqrt{3}}{\sqrt{25}}\]

OpenStudy (lgbasallote):

@purplec16 next time show some solution so our askers can learn eh? ;)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large \sqrt{3 \over 25} = {\sqrt{3} \over \sqrt{25}} = {\sqrt3 \over 5 }}\)

OpenStudy (lgbasallote):

oh you did..next time post the solution before answer :p lol

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome!